# When Simultaneous Localization and Mapping Meets Wireless Communications: A Survey

### v1

@@ IV-B3 Visual Range Odometry-based SLAM
Significant efforts have been made to achieve accurate real-time navigation using SLAM and pose estimation in environments and applications where high accuracy can practically serve an application scope. An indoor wayfinding system based on geometric features aided graph SLAM for the visually impaired has been presented in [144]. The authors of this study addressed the challenge of indoor navigation within the framework of a robotic navigation aid, proposing a client-server architecture for real-time wayfinding. Producing both range data and intensity images at rates of up to 50 FPS, the SR4000 is well-suited for SLAM applications due to its compact dimensions (50 × 48 × 65 mm³) and configurable modulation frequency, allowing multiple cameras to operate simultaneously without interference. In terms of architecture, the client device captures data from the camera and transmits it to the server via WiFi while also interacting with the user through a speech interface. The server, in turn, performs SLAM, localizes the user within a floor plan, plans the optimal path, and sends navigational commands and points-of-interest messages back to the client. To estimate the camera’s egomotion, the system employed the visual range odometry (VRO) algorithm [86], which extracts visual features from the current intensity image and matches them with those in the previous frame. The SIFT feature descriptors [77] facilitate feature matching, while siftGPU [112] accelerates feature extraction and matching, achieving a runtime of approximately 30 milliseconds for VRO. The SLAM approach follows a two-step process: first, it extracts the floor plane from the camera’s point cloud data and integrates it as a landmark in a 3-D graph SLAM process, reducing pitch, roll, and Z-axis errors. Next, it projects the point data onto the extracted floor plane to detect wall lines, incorporating them into a 2D graph SLAM (in the OXY plane) to minimize yaw, X, and Y errors. By leveraging geometric features-namely, floor planes and wall lines-this two-step SLAM method effectively reduced 6-DOF pose estimation errors.
@@ IV-B4 RSSI-based Multi-agent Localization
IV-C DNN-based Wireless SLAM
@@
IV-C1 Deep Coupling SLAM
@@
IV-C2 Deep Semantic and Spatial Joint Perception
@@
IV-D Radio Maps Extraction Algorithm via mmWave Sensing
@@
Figure 6: EZMAP: a high-accuracy, low-cost floor plan construction system that fuses RF tracking with inertial sensing, based on [155]
@@
The study in [76] explored the challenges posed by next-generation wireless communication, where user devices autonomously scan their surroundings by steering narrow antenna beams and collecting echoes from objects and walls. The authors developed radio-SLAM algorithms inspired by image signal processing techniques, leveraging mobile THz radar for environment mapping and device localization. Their approach treats the channel transfer function as an angle-delay matrix, where back-scattered channel impulse responses (CIRs) provide rich spatial information. Central to this study is a pose estimation algorithm operating in the radar coordinate system, incorporating an EKM for accurate localization. The measurements used for estimation include horizontal and vertical position shifts, as well as rotational deviations relative to previous poses. A notable innovation in the same study is the Fourier-Mellin-based algorithm, a fast Fourier transform (FFT)-based technique traditionally used for image registration by optimizing matches in the frequency domain [76]. The study applied this algorithm to consecutive radar frames in polar coordinates, treating them as ‘radio images‘ of the environment. By decoupling translation and rotation effects, the method improves relative pose estimation between two consecutive frames. The authors compared the Fourier-Mellin approach with the Laser Scan Matching algorithm, demonstrating superior performance in radio-based localization.
@@
$f_{D}=\frac{2f}{c}(\mathbf{v}+\boldsymbol{\omega}\times\mathbf{p})^{T}\cdot\mathbf{n},$
@@
IV-E WiFi Fingerprinting-based Localization
@@
IV-F Device-based Sensing and Localization
@@
The study in [140] considers a mobile mmWave device sensing application, exploiting the small wavelength of mmWave signals to deploy large antenna arrays at both the transmitter and receiver. 5G mmWave communications are deemed to be inherently facilitating SLAM-enabled UE, capitalizing on larger bandwidths and antenna array augmentation as previously mentioned [134]. Compared to sub-6 GHz communication, mmWave technology offers superior delay and angle measurements, making it well-suited for SLAM applications These large arrays provide improved angular resolution, enabling precise environmental mapping. The transmitter scans the entire angular space, while the receiver captures echo signals reflected from scatterers in the environment. The locations of these scatterers are then extracted from the received signals, allowing the system to construct a spatial map of the surroundings. Extending the discussion to 6D pose estimation, the authors of [109] explore how UE can determine both position and orientation. They highlight the ability to integrate the huge antenna arrays into user devices while maintaining a small physical footprint, unlocking new possibilities for UE-based environmental awareness.
@@
IV-G Reconfigurable Intelligent Surfaces (RIS)-aided SLAM
@@
V Challenges and Future Directions
The deployment of SLAM-enabled networked autonomous systems in real-world environments reveals a set of fundamental challenges that stem from the mismatch between idealized modeling assumptions and the complex dynamics of practical operation. Among these challenges, environment dynamics play a central role, since both the physical scene and the wireless propagation conditions evolve continuously over time. In contrast to classical SLAM formulations that assume static landmarks and slowly varying sensor parameters, real world scenarios are characterized by moving objects, time varying occlusions, and rapidly changing geometry. As such, motion has a direct impact on reliability and estimation consistency [10]. V-SLAM pipelines can experience unstable feature tracking and erroneous data association under illumination changes, motion blur, and partial occlusions, leading to map degradation and pose drift [106, 82]. Moreover, wireless-based localization and SLAM methods operating at mmWave and higher frequencies are highly sensitive to propagation variations, where blockage, Doppler shifts, and the intermittent appearance of dominant multipath components cause abrupt fluctuations in received signal strength, angle estimates, and CSI [95, 120]. As a result, when visual and wireless modalities are jointly considered, these dynamics further challenge system robustness, since the two sensing modalities respond differently to environmental changes and operate at different temporal resolutions, leading to cross modal inconsistencies and outdated representations [84]. As such, maintaining accurate localization and mapping in the presence of dynamic geometry and nonstationary sensing conditions remains an open challenge, motivating the development of SLAM frameworks that explicitly account for temporal variability and partial observability rather than relying on static environment assumptions.
@@ V Challenges and Future Directions
Finally, as SLAM-enabled networked autonomous systems increasingly rely on connectivity, distributed processing, and data sharing, attributes such as security and privacy emerge as additional challenges derived from the shift toward networked perception. In particular, the exchange of localization and mapping information among multiple agents or with edge and cloud infrastructures enlarges the system attack surface, and therefore exposes SLAM pipelines to cyber and physical threats that can directly affect estimation reliability and operational safety. Recent studies have demonstrated that adversarial manipulation of sensory inputs or communication links can corrupt localization and mapping results, which, in turn, can lead to unsafe or misleading behavior in autonomous platforms [70, 94]. Vision-based SLAM systems are susceptible to visual spoofing and adversarial perturbations, whereas wireless-assisted SLAM frameworks may be affected by jamming, spoofing, or falsified CSI. The aforementioned can cause a degradation in both localization accuracy and communication performance [56]. Beyond integrity attacks, SLAM data inherently encodes sensitive information about environments and agent trajectories, including spatial layouts, mobility patterns, and potentially identifiable visual content, and thus raises significant privacy concerns when such data is stored processed or shared beyond the local platform [39]. These concerns are amplified in multimodal and networked settings. The fusion of visual and wireless data enables richer inference capabilities, at the cost, however, of increasing the potential for unintended information leakage and cross-domain attacks. Thus, security and privacy must be considered core challenges in SLAM-enabled networked autonomous systems, motivating the development of frameworks that incorporate robust estimation under adversarial conditions, secure communication mechanisms [89], and privacy-aware data representations as integral design components rather than auxiliary protections.
@@
Based on the challenges discussed above, future research on SLAM-enabled networked autonomous systems is expected to further explore approaches that address environment dynamics, latency, energy efficiency, and security in an integrated manner. Since these challenges are tightly coupled at the system level, improving a single component is often sub-optimal, thereby motivating more holistic design principles that jointly consider sensing, communication, and computation. Recent advances in wireless communications, edge intelligence, and multimodal perception have led to increased interest in frameworks where SLAM and communication subsystems interact more closely, particularly in dynamic and networked environments. Rather than treating localization and mapping as independent of wireless operation, several works have investigated whether combining visual observations with radio measurements can improve robustness under fast-changing geometry and intermittent sensing conditions [84]. In vehicular and mobile robotic scenarios, for example, visual data acquired from low-cost sensors such as monocular cameras can be fused with wireless indicators including received signal strength or CSI to jointly estimate agent pose and motion-related states. By exploiting temporal correlations across multimodal data streams, such approaches may enable short-term prediction of both mobility and wireless link dynamics, which is particularly relevant in environments affected by frequent occlusions and non-line-of-sight transitions. In addition, these predictive capabilities can support proactive communication functions, such as beam selection or handover decisions, while potentially reducing the need for frequent channel probing, and thus alleviating latency and energy constraints. Conversely, wireless measurements can provide complementary information when visual perception is degraded, offering an additional sensing modality that may help mitigate failures caused by adverse lighting or occlusions. Finally, recent studies on joint communication and sensing architectures suggest that such bidirectional interaction between perception and communication layers could form a foundation for SLAM-aware communication systems and communication-aware SLAM frameworks, especially in beyond-5G and 6G networked autonomy settings [127]. Overall, these directions highlight a broader shift toward tightly integrated, predictive, and resource-aware SLAM architectures, where robustness, efficiency, and trustworthiness must be addressed jointly in order to enable reliable operation in complex real-world environments.
VI Conclusions
@@
This survey has explored the emerging nexus between SLAM and wireless communications, focusing on their growing interdependence in the era of mmWave/GHz communications, 5G, and beyond. The study highlighted how RF and visible light spectra camera modalities can serve as the basis for networked autonomous systems’ to converge towards a unified localization, mapping, communication, and control framework. Through an integrative taxonomy incorporating the foundations of SLAM in networked autonomous systems, SLAM hetero-modal techniques as well as mathematical methods, and wireless communication–enabled perception, the study was oriented towards the aforementioned convergence.
@@
A key outcome of the information synthesis that was endeavored is the identification of the true bidirectional relationship between wireless systems and SLAM. On one hand, wireless communications, through RF, WiFi, and mmWave signals, enable more accurate localization and NLOS situational awareness, especially under occluded visual conditions. On the other hand, computer vision and SLAM methodologies, including visual odometry, feature-based tracking for translational and rotational velocity retrieval, and feature-based landmark reconstruction, offer new tools for geometric modeling of communication channels, beam prediction, and mobility-aware network modeling. This reciprocity demonstrates how perception and connectivity can evolve in parallel within a shared estimation and motion planning loop.
@@
At an estimation-theoretical level, the survey revisited probabilistic modeling approaches in localization such as Bayesian filtering, belief-propagation-based state estimation whilst more sophisticated approaches such as vector-field SLAM built on a mathematical background that stood out from the crowd given an observed trend that equalizes SLAM with extended Kalman filtering. From the technological perspective of wireless communication, the discussion included DL–based localization, radio map construction, RIS-aided MetaSLAM, and multi-agent active estimation frameworks, illustrating the deep convergence of communication, sensing, and control. The study also highlighted some emerging trends towards ISAC, where wireless infrastructures act as both communication media and spatial sensors. To this end, passive SLAM was identified as a natural evolution of state observation in the context of networked autonomy, where perception is inherently maintained by rich visual and wireless signals without deliberately orienting the networked autonomous vehicle towards the direction of information gain maximization or estimation uncertainty gradient descent. Active SLAM, on the other hand, utilizes the system’s control input vector to actively reduce uncertainty for a computer-vision-driven such as UAV cinematography or communications-driven such as beam tracking perception goal.
@@
The findings of this study indicate that joint SLAM–wireless systems design will form a cornerstone for future networked autonomous systems, enabling safe navigation, robust state feedback control, environment reconstruction, and real-time network optimization. Whilst early research demonstrates promising results, fully integrated solutions are still in their infancy. Advancements in hetero-modal fusion, distributed estimation information exchange among autonomous agents, vision-language models in robotics, and edge-powered computation are deemed important in order to achieve the next generation of visual perception-aware communications and communications-aware visual perception. Thus, the nexus of SLAM and wireless communications is a foundational paradigm for 6G-enabled intelligent robotic and autonomous vehicle systems, capable of advanced reasoning on space and time through adaptive, unobstructed wireless links.

### v2

@@ IV-B3 Visual Range Odometry-based SLAM
Significant efforts have been made to achieve accurate real-time navigation using SLAM and pose estimation in environments and applications where high accuracy can practically serve an application scope. An indoor wayfinding system based on geometric features-aided graph SLAM for the visually impaired has been presented in [147]. The authors of this study addressed the challenge of indoor navigation within the framework of a robotic navigation aid, proposing a client-server architecture for real-time wayfinding. Producing both range data and intensity images at rates of up to 50 FPS, the SR4000 is well-suited for SLAM applications due to its compact dimensions (50 × 48 × 65 mm³) and configurable modulation frequency, allowing multiple cameras to operate simultaneously without interference. In terms of architecture, the client device captures data from the camera and transmits it to the server via WiFi while also interacting with the user through a speech interface. The server, in turn, performs SLAM, localizes the user within a floor plan, plans the optimal path, and sends navigational commands and points-of-interest messages back to the client. To estimate the camera’s egomotion, the system employed the visual range odometry (VRO) algorithm [88], which extracts visual features from the current intensity image and matches them with those in the previous frame. The SIFT feature descriptors [79] facilitate feature matching, while siftGPU [114] accelerates feature extraction and matching, achieving a runtime of approximately 30 milliseconds for VRO. The SLAM approach follows a two-step process: first, it extracts the floor plane from the camera’s point cloud data and integrates it as a landmark in a 3-D graph SLAM process, reducing pitch, roll, and Z-axis errors. Next, it projects the point data onto the extracted floor plane to detect wall lines, incorporating them into a 2D graph SLAM (in the OXY plane) to minimize yaw, X, and Y errors. By leveraging geometric features-namely, floor planes and wall lines-this two-step SLAM method effectively reduced 6-DOF pose estimation errors.
@@ IV-B4 RSSI-based Multi-agent Localization
Table IV presents an overview of the discussed SLAM modality-based techniques, highlighting their key characteristics.
TABLE IV: Overview of SLAM Modality-Based Techniques and Core Characteristics
SLAM Modality Key Characteristics Core Sensing Technologies & Algorithms Representative Studies
Inertial SLAM 3-D collaborative occupancy grid-based inertial SLAM; switching floor event matching; effective motion estimation under limited external metric references. MIMU; pedestrian navigation systems (PNS); visual-inertial navigation systems (VINS). [69, 95, 56]
LiDAR-based SLAM Point cloud mapping extendable by GNSS and inertial sensing; improving indirectly the accuracy of pose optimization; robust long-term localization in structured environments within area graphs (AGLoc); augmenting wireless-enabled frameworks utilizing WiFi APs. LRF; LiDAR-extended sensor fusion; AGLoc; LiDAR & WiFi & camera-based fusion; ICP. [43, 12, 131, 52, 155, 49]
Visual Range Odometry-based SLAM Visual range odometry using feature extraction and matching; GPU-accelerated SIFT descriptors; real-time pose estimation and mapping. 3-D cameras; SIFT feature descriptors; GPU acceleration; visual odometry. [147, 88, 79, 114]
Multi-agent RSSI-based Localization Beacon-based cooperative localization using RSSI measurements; Bayesian inference; multi-robot coordination. RSSI sensing; wireless beacons; multi-robot systems; cooperative localization frameworks. [55]
Lessons Learned:
  1. 1.
    Inertial SLAM can serve as an alternative to V-SLAM in indoor environments, enabling 3-D occupancy grid reconstruction and motion estimation, including handling of switching floor events. LiDAR-based SLAM naturally provides dense 3-D point cloud representations, which, combined with algorithms such as DBSCAN and ICP, enable robust tracking of structural entities. Often integrated within sensor fusion frameworks, LiDAR combined with visual, inertial, and wireless systems can further enhance performance. Visual range odometry-based SLAM, particularly when GPU-accelerated, can support accurate real-time pose estimation and mapping. In view of cooperative localization, multi-agent systems can provide state estimates using RSSI measurements and relying on Bayesian estimation frameworks.
V Wireless- and 6G-Enabled SLAM
Wireless communications in the era of beyond-5G are considered a key enabler of AI and especially ML/DL frameworks that aim to perform in real-time. Whilst the high transmission data rates and the effective retrieval of deep embeddings sent over wireless links provide the capabilities to support autonomous systems’ applications, properties at the physical layer, through the utilized wider spectrum of high frequencies, can be particularly important for localization and mapping. In this section, radio waves are considered as SLAM-related input features that form high-dimensional spaces, resulting in the learning of representations in a manner analogous to computer vision ML/DL embeddings.
V-A DNN-based Wireless SLAM
@@
V-A1 Deep Coupling SLAM
@@
V-A2 Deep Semantic and Spatial Joint Perception
@@
V-B Radio Maps Extraction Algorithm via mmWave Sensing
@@
Figure 11: EZMAP: a high-accuracy, low-cost floor plan construction system that fuses RF tracking with inertial sensing, based on [158].
@@
The study in [78] explored the challenges posed by next-generation wireless communication, where user devices autonomously scan their surroundings by steering narrow antenna beams and collecting echoes from objects and walls. The authors developed radio-SLAM algorithms inspired by image signal processing techniques, leveraging mobile THz radar for environment mapping and device localization. Their approach treats the channel transfer function as an angle-delay matrix, where back-scattered channel impulse responses (CIRs) provide rich spatial information. Central to this study is a pose estimation algorithm operating in the radar coordinate system, incorporating an EKF for accurate localization. The measurements used for estimation include horizontal and vertical position shifts, as well as rotational deviations relative to previous poses. A notable innovation in the same study is the Fourier-Mellin-based algorithm, an FFT-based technique traditionally used for image registration by optimizing matches in the frequency domain [78]. The study applied this algorithm to consecutive radar frames in polar coordinates, treating them as ‘radio images‘ of the environment. By decoupling translation and rotation effects, the method improves relative pose estimation between two consecutive frames. The authors compared the Fourier-Mellin approach with the Laser Scan Matching algorithm, demonstrating superior performance in radio-based localization.
@@
$f_{D}=\frac{2f}{c}(\mathbf{v}+\boldsymbol{\omega}\times\mathbf{p})^{T}\cdot\mathbf{n},$ (43)
@@
V-C WiFi Fingerprinting-based Localization
@@
V-D Device-based Sensing and Localization
@@
The study in [142] considers a mobile mmWave device sensing application, exploiting the small wavelength of mmWave signals to deploy large antenna arrays at both the transmitter and receiver. 5G mmWave communications are deemed to be inherently facilitating SLAM-enabled UE, capitalizing on larger bandwidths and antenna array augmentation as previously mentioned [136]. Compared to sub-6 GHz communication, mmWave technology offers superior delay and angle measurements, making it well-suited for SLAM applications. These large arrays provide improved angular resolution, enabling precise environmental mapping. The transmitter scans the entire angular space, while the receiver captures echo signals reflected from scatterers in the environment. The locations of these scatterers are then extracted from the received signals, allowing the system to construct a spatial map of the surroundings. Extending the discussion to 6D pose estimation, the authors of [111] explore how UE can determine both position and orientation. They highlight the ability to integrate the huge antenna arrays into user devices while maintaining a small physical footprint, unlocking new possibilities for UE-based environmental awareness.
@@
V-E Reconfigurable Intelligent Surfaces (RIS)-aided SLAM
@@
VI Challenges and Future Directions
VI-A Challenges
The deployment of SLAM-enabled networked autonomous systems in real-world environments reveals a set of fundamental challenges that stem from the mismatch between idealized modeling assumptions and the complex dynamics of practical operation.
VI-A1 Illumination, Propagation, and Occlusion Effects
Among these challenges, environment dynamics play a central role, since both the physical scene and the wireless propagation conditions evolve continuously over time. In comparison with classical SLAM formulations that assume static landmarks and slowly varying sensor parameters, real world scenarios are characterized by moving objects, time varying occlusions, and rapidly changing geometry. As such, motion could directly affect reliability and estimation consistency [11]. V-SLAM pipelines can experience unstable feature tracking and erroneous data association under illumination changes, motion blur, and partial occlusions, leading to map degradation and pose drift [108, 84]. Moreover, wireless-based localization and SLAM methods operating at mmWave and higher frequencies are highly sensitive to propagation variations, where blockage, Doppler shifts, and the intermittent appearance of dominant multipath components cause abrupt fluctuations in received signal strength, angle estimates, and CSI [97, 122]. As a result, when visual and wireless modalities are jointly considered, these dynamics further challenge system robustness, since the two sensing modalities respond differently to environmental changes and operate at different temporal resolutions, leading to cross modal inconsistencies and outdated representations [86]. As such, maintaining accurate localization and mapping in the presence of dynamic geometry and nonstationary sensing conditions remains an open challenge, motivating the development of SLAM frameworks that account for temporal variability and partial observability rather than relying on static environment assumptions.
VI-A2 Latency Constraints
@@ VI-A4 Security and Privacy
Finally, as SLAM-enabled networked autonomous systems increasingly rely on connectivity, distributed processing, and data sharing, attributes such as security and privacy emerge as additional challenges derived from the shift toward networked perception. In particular, the exchange of localization and mapping information among multiple agents or with edge and cloud infrastructures enlarges the system attack surface, and therefore exposes SLAM pipelines to cyber and physical threats that can directly affect estimation reliability and operational safety. Recent studies have shown that adversarial manipulation of sensory inputs or communication links can corrupt localization and mapping results, which, in turn, can lead to unsafe or misleading behavior in autonomous platforms [72, 96]. Vision-based SLAM systems are susceptible to visual spoofing and adversarial perturbations, whereas wireless-assisted SLAM frameworks may be affected by jamming, spoofing, or falsified CSI. The aforementioned can cause a degradation in both localization accuracy and communication performance [58]. Beyond integrity attacks, SLAM data inherently encodes sensitive information about environments and agent trajectories, including spatial layouts, mobility patterns, and potentially identifiable visual content, and thus raises significant privacy concerns when such data is stored, processed, or shared beyond the local platform [41]. These concerns are amplified in multimodal and networked settings. The fusion of visual and wireless data enables richer inference capabilities, at the cost, however, of increasing the potential for unintended information leakage and cross-domain attacks. Thus, security and privacy can be considered core challenges in SLAM-enabled networked autonomous systems, motivating the development of frameworks that incorporate robust estimation under adversarial conditions, secure communication mechanisms [92], and privacy-aware data representations as integral design components rather than auxiliary protections.
VI-B Future Directions
@@
Based on the challenges discussed above, future research on SLAM-enabled networked autonomous systems is expected to further explore approaches that address environment dynamics, latency, energy efficiency, and security in an integrated manner. Since these challenges are tightly coupled at the system level, improving a single component is often sub-optimal, thereby motivating more holistic design principles that jointly consider sensing, communication, and computation. Recent advances in wireless communications, edge intelligence, and multimodal perception have led to increased interest in frameworks where SLAM and communication subsystems interact more closely, particularly in dynamic and networked environments. Rather than treating localization and mapping as independent of wireless operation, several works have investigated whether combining visual observations with radio measurements can improve robustness under fast-changing geometry and intermittent sensing conditions [86]. In vehicular and mobile robotic scenarios, for example, visual data acquired from low-cost sensors such as monocular cameras can be fused with wireless indicators including received signal strength or CSI to jointly estimate agent pose and motion-related states. By exploiting temporal correlations across multimodal data streams, such approaches may suggest short-term prediction of both mobility and wireless link dynamics, which is particularly relevant in environments affected by frequent occlusions and NLOS transitions. In addition, these predictive capabilities can support proactive communication functions, such as beam selection or handover decisions, while potentially reducing the need for frequent channel probing, and thus alleviating latency and energy constraints. Conversely, wireless measurements can provide complementary information when visual perception is degraded, offering an additional sensing modality that may help mitigate failures caused by adverse lighting or occlusions. Finally, recent studies on joint communication and sensing architectures suggest that such bidirectional interaction between perception and communication layers could form a foundation for SLAM-aware communication systems and communication-aware SLAM frameworks, especially in beyond-5G and 6G networked autonomy settings [129]. Overall, these directions highlight a broader shift toward tightly integrated, predictive, and resource-aware SLAM architectures, where robustness, efficiency, and trustworthiness addressed jointly may warrant further research in real-world environments.
VII Conclusions
@@
This survey has examined elements of the emerging intersection between SLAM and wireless communications, focusing on their growing interdependence in the era of mmWave/GHz communications, 5G, and beyond. The study highlighted how RF and visible light spectra camera modalities can serve as the basis for networked autonomous systems’ to converge towards a unified localization, mapping, communication, and control framework. Through a framework integrating the foundations of SLAM in networked autonomous systems, SLAM hetero-modal techniques, mathematical methods, and wireless communication–enabled perception, the study was oriented towards the aforementioned convergence.
@@
A key outcome of the information synthesis that was undertaken is the identification of the bidirectional relationship between wireless systems and SLAM. On one hand, wireless communications, through RF, WiFi, and mmWave signals, have the potential for accurate localization and NLOS situational awareness, especially under occluded visual conditions. On the other hand, computer vision and SLAM methodologies, including visual odometry, feature-based tracking for translational and rotational velocity retrieval, and feature-based landmark reconstruction, providing new tools for geometric modeling of communication channels, beam prediction, and mobility-aware network modeling. This reciprocity shows how perception and connectivity can evolve in parallel within a shared estimation and motion planning loop.
@@
At an estimation-theoretical level, the survey revisited probabilistic modeling approaches in localization such as Bayesian filtering, belief-propagation-based state estimation whilst more sophisticated approaches such as vector-field SLAM built on a mathematical background that stood out from the crowd given an observed trend that equalizes SLAM with extended Kalman filtering. From the technological perspective of wireless communication, the discussion included DL–based localization, radio map construction, RIS-aided MetaSLAM, and multi-agent active estimation frameworks, suggesting deep convergence of communication, sensing, and control. The study also highlighted some emerging trends towards ISAC, where wireless infrastructures can act as both communication media and spatial sensors. To this end, passive SLAM was identified as a natural evolution of state observation in the context of networked autonomy, where perception is inherently maintained by rich visual and wireless signals without deliberately orienting the networked autonomous vehicle towards the direction of increasing information gain or estimation uncertainty gradient decline. Active SLAM, on the other hand, can utilize the system’s control input vector to actively reduce uncertainty for a computer-vision-driven such as UAV cinematography or communications-driven such as beam tracking perception goal.
@@
The findings of this study indicate that joint SLAM–wireless systems design has the potential to form a cornerstone for future networked autonomous systems, facilitating safe navigation, robust state feedback control, environment reconstruction, and real-time network optimization. Whilst early research shows promising results, fully integrated solutions appear to be in their infancy. Advancements in hetero-modal fusion, distributed estimation information exchange among autonomous agents, vision-language models in robotics, and edge-powered computation are important for achieving the next generation of visual perception-aware communications and communications-aware visual perception. Thus, the intersection of SLAM and wireless communications forms a foundational paradigm for 6G-enabled intelligent robotic and autonomous vehicle systems, supporting advanced reasoning on space and time through adaptive, unobstructed wireless links.